# Non-radial oscillations of hadronic neutron stars, quark stars, and hybrid stars : Estimation of $f$, $p$, and $g$ modes

### v1

@@ Abstract
The composition and equation of state (EoS) of dense matter relevant to compact stars are quite inconclusive. However, certain observational constraints on the structural properties of compact stars help us constrain the EoS to a fair extent. Moreover, gravitational asteroseismology gives us a notion of the interior of the compact star viz., its composition and EoS. The next generation gravitational wave (GW) detectors are likely to detect several oscillation mode frequencies of the GWs, specially the fundamental ($f$) mode frequency. In this work we compute the $f$ and the first pressure ($p_{1}$) mode frequencies ($f_{f}$ and $f_{p1}$, respectively) with different compositions viz., hadronic matter, quark matter, and hybrid star (HS) matter. For HSs, we also study the gravity ($g$) mode frequency ($f_{g}$) arising due to the discontinuity in density. For each phase we also study the correlation between the oscillation frequencies of the 1.4 $M_{\odot}$ and 2.01 $M_{\odot}$ compact stars with the other different properties. We find that the different possible composition of compact stars substantially affects the different oscillation mode frequencies. However, the mass-scaled angular $f$ mode frequency ($\omega_{f}M$) varies universally with compactness ($C$) for all hadronic, quark and hybrid stars. The $f$ mode frequency ($f_{f_{1.4}}$) of the canonical 1.4 $M_{\odot}$ compact star, obtained with the different composition, is quite correlated with the canonical radius ($R_{1.4}$) and tidal deformability ($\Lambda_{1.4}$) while $f_{p_{1.4}}$ is well correlated with the slope parameter ($L$) of the symmetry energy. We also show that $f_{g_{1.4}}$ of the HSs varies almost linearly with $\Lambda_{1.4}$. Should $g$ modes be detected, they could not only support the existence of HSs, but the magnitude of $f_{g}$ could be useful to understand the strength of quark repulsion in HSs.
@@ I Introduction
Several extreme conditions of density ($\rho$ = 5 - 10 times nuclear matter density), compactness ($C$ = $M/R$), and etc. related to the neutron star (NS) environment make them one of the most fascinating celestial objects to study. However, till date a lot of uncertainties still pertain to their equation of state (EoS) which is developed on the basis of theoretical modeling of NS matter. These uncertainties are generally attributed to the fact that the composition, properties and interaction of matter at such high density are largely inconclusive. Over several decades theoretical speculations suggest that NS cores are dense enough to sustain the existence of exotic matter and even signatures of hadron-quark phase transition thereby forming hybrid stars (HSs). In order to achieve phase transition, mechanisms like Gibbs and Maxwell constructions Maruyama et al. (2008); Orsaria et al. (2019), hadron-quark crossover Qin et al. (2023); Constantinou et al. (2021); Sotani and Kojo (2023) and constant speed of sound parameterization Sun and Wen (2023); Laskos-Patkos and Moustakidis (2023) etc. are often adopted in literature. The Bodmer-Witten conjecture suggests that strange quark matter (SQM), being more stable than pure nucleonic matter, is the true ground state of the hadronic matter Farhi and Jaffe (1984); Torres and Menezes (2013); Olinto (1987). This implies that the entire compact star may be composed of SQM and thereby strange quark stars (SQSs) may also exist. At very high density, QM can also give rise to the formation of color-flavor-locked (CFL) QM Rajagopal and Wilczek (2001); Alford and Reddy (2003); Lugones and Horvath (2002); Vásquez Flores and Lugones (2017). Consequently, based on the composition and interactions, we obtain the EoS of hadronic neutron stars (HNSs), SQSs, and HSs. Each class of EoS has its own uncertainty revolving around the consideration of composition and interactions. With the recent developments in the observational sector, the EoS has been constrained to a certain extent. The most important of such observational constraints include the detection of gravitational wave GW170817 by the LIGO/Vigro Collaboration Abbott et al. (2018) and the simultaneous measurement of mass and radius of HESS J1731-347 V. Doroshenko, V. Suleimanov, G. Pühlhofer, and A. Santangelo (2022) and those of PSR J0030+0451 Riley et al. (2019); Miller et al. (2019) and PSR J0740+6620 Fonseca et al. (2021); Miller et al. (2021); Riley et al. (2021) by NICER experiment.
@@
With the recent advancements of the upcoming gravitational wave (GW) detectors like the the LIGO O4 run, Einstein Telescope, and the Cosmic Explorer, it can be expected that the oscillation mode frequencies of the GWs can be detected, which will provide deeper insight to the interior of the compact stars. Such oscillation spectra of GWs consist of different modes viz., fundamental ($f$), pressure ($p$), rotational ($r$), space-time ($w$) and gravity ($g$) modes Kokkotas and Schmidt (1999). The restoring force that re-establishes the equilibrium is different for each mode. A lot of studies have been done on these different oscillation modes Sotani et al. (2011); Flores and Lugones (2014); Andersson and Pnigouras (2019); Zhao and Lattimer (2022); Kumar et al. (2024a); Ranea-Sandoval et al. (2018); Vásquez Flores et al. (2019); Pradhan et al. (2022); Jyothilakshmi et al. (2024); Das et al. (2021); Hong et al. (2023); Shirke et al. (2024); Jaiswal and Chatterjee (2021); Constantinou et al. (2021); Orsaria et al. (2019); Kumar et al. (2023a); Zheng et al. (2023); Sen and Guha (2024); Wen et al. (2019); Sotani (2021) based on different composition of the star. In case of both the $f$ and $p$ modes, the equilibrium is restored by fluid pressure of the star and for the $g$ mode, gravity acts as the restoring force. The efficiency of the upcoming detectors is expected to aid the branch of GW asteroseismology, especially in terms of the most prominent $f$ mode frequency ($f_{f}$ = 1-3 kHz). Thus in the present and recent future, it is important to compute the oscillation mode frequencies with the different EoS. Several works dedicated to this topic suggest strong correlation between the $f_{f}$ and compactness Andersson and Kokkotas (1998), moment of inertia Lau et al. (2010) and static tidal polarizability Chan et al. (2014); Sotani and Kumar (2021) for NSs, QSs and HSs Zhao and Lattimer (2022), leading to the emergence of the universal relations between the $f_{f}$ and such properties of compact stars. $p$ mode frequency ($f_{p1}$), on the other hand, are weakly correlated with such properties of NSs Andersson and Kokkotas (1998). Moreover, they are largely affected by the EoS of the crust, unlike the $f_{f}$, especially at low density Vásquez Flores et al. (2017); Pradhan and Chatterjee (2021); Kunjipurayil et al. (2022). The $g$ mode and its corresponding frequency $f_{g}$ arise due to the discontinuity in density of the EoS of the NS which may be attributed to the phenomena like first order phase transition and thus it can be described by HS EoS Ranea-Sandoval et al. (2019); Flores and Lugones (2014); Sotani et al. (2011); Ranea-Sandoval et al. (2018); Tonetto and Lugones (2020); Wei et al. (2020); Rodriguez et al. (2021); Kumar et al. (2023b); Zhao et al. (2022); Tran et al. (2023); Zhang et al. (2024).
@@
In the present work we study the effects of different composition of compact star matter on the non-radial oscillation frequencies of the $f$ and $p$ modes of the HNSs and SQSs with different models. For HSs we also calculate the $g$ mode frequency $f_{g}$. The first development of the theory of non-radial oscillation of NS in the framework of general relativity (GR) is done in Thorne and Campolattaro (1967) while Lindblom and Detweiler (1983) first integrated the numerical solution of the NS oscillation. Later the metric perturbations were neglected with the help of Cowling approximations Cowling (1941) in order to obtain a simplified theory Sotani et al. (2011). The calculation of the oscillation frequencies in full GR conditions that include the spacetime oscillation also shows the emergence of another mode viz., the $w$ mode. Therefore, many recent works have calculated the oscillation frequencies in the framework of the total GR theory Shirke et al. (2024); Mondal and Bagchi (2023); Zhao and Lattimer (2022); Kunjipurayil et al. (2022); Pradhan et al. (2022); Flores et al. (2024); Tsui and Leung (2005). The results of the calculation of the oscillation frequencies using Cowling approximation and total GR calculations can differ up to 30% for the $f_{f}$, and $\sim$15for $f_{p1}$ Kunjipurayil et al. (2022). However, the overestimation with Cowling approximation can be less than 20% from those calculated by considering the full GR treatment Pradhan et al. (2022); Yoshida and Kojima (1997). Also, Tonetto and Lugones (2020); Zhao et al. (2022) showed that the Cowling approximation is reasonably good for calculation of $f_{g}$. Overall, the gross qualitative results remain unaffected with or without Cowling approximation. Therefore, it is still widely adopted in recent literature Kumar et al. (2023b); Jyothilakshmi et al. (2024); Arbañil et al. (2023); Thakur et al. (2024); Thapa et al. (2023) and in the present work we also consider the Cowling approximation in order to provide the estimates of different oscillation frequencies. For the purpose to explore the uncertainty due to the EoS, we consider the KIDS (Korea-IBS-Daegu-SKKU) functional Gil et al. (2021); Gil and Hyun (2021); Gil et al. (2022), and the relativistic mean field (RMF) models GM1 Glendenning and Moszkowski (1991) and DD2 Typel et al. (2010) hadronic models. For SQSs we consider the vector Bag (v-Bag) model Torres and Menezes (2013); Lopes et al. (2022), the Nambu-Jona-Lasino (NJL) model with vector interaction Lopes and Menezes (2021); Masuda et al. (2013), and the CFL model based on the MIT Bag model Lugones and Horvath (2002). For HSs, we follow Maxwell construction Kumar et al. (2023c); Qin et al. (2023); Laskos-Patkos et al. (2024) with the KIDS model to describe the hadronic phase and the v-Bag model for the quark phase. We intend to study how the oscillation properties vary for the three different types of stars viz. HNSs, SQSs and HSs based on the composition. We mention in this context that the KIDS model is not considered before in literature to study the oscillation properties of HNSs and HSs.
@@
We organize this paper in the following manner. In the next Sec. II we discuss the salient features of the different models considered in the present work on the basis of different composition. Section III is devoted to the description of hadron-quark phase transition. We also provide the essentials of the mechanism to obtain the structural properties and different modes of non-radial oscillation of compact star in Sec. IV. We display our results with corresponding discussion in Sec. V. We finally conclude in Sec. VI.
@@ II Models for hadronic and quark matters
We consider the KIDS functional Gil et al. (2021); Gil and Hyun (2021); Gil et al. (2022) as the hadronic model to compute the structural properties and the oscillations frequencies of HNSs. We compare the results obtained with the KIDS model, with those of the two RMF models viz., the GM1 Glendenning and Moszkowski (1991) and the DD2 Typel et al. (2010) models. To obtain the structural properties and the oscillations frequencies of the QSs we choose to work with different quark matter models viz., the vector Bag (v-Bag) model Torres and Menezes (2013); Lopes et al. (2022), the Nambu-Jona-Lasino (NJL) model with vector interaction Lopes and Menezes (2021); Masuda et al. (2013), and the color-flavor-locked (CFL) model based on the MIT Bag model Lugones and Horvath (2002). The salient features of each of the hadronic and quark models, considered in the present work, are briefly addressed below.

### v2

# Non-radial oscillations of hadronic neutron stars, quark stars, and hybrid stars : Calculation of $f$, $p$, and $g$ mode frequencies

@@ Abstract
The composition and equation of state (EoS) of dense matter relevant to compact stars are quite inconclusive. However, certain observational constraints on the structural properties of compact stars help us constrain the EoS to a fair extent. Moreover, gravitational asteroseismology gives a notion of the composition and EoS of compact stars. The next generation gravitational wave (GW) detectors are likely to detect several oscillation mode frequencies of the GWs. In this work we compute the fundamental ($f$) and the first pressure ($p_{1}$) mode frequencies ($f_{f}$ and $f_{p1}$, respectively) with different compositions viz., hadronic, quark, and hybrid star (HS) matter. For HSs, we also study the gravity ($g$) mode frequency ($f_{g}$). For each phase we also study the correlation between the oscillation frequencies of 1.4 $M_{\odot}$ and 2.01 $M_{\odot}$ compact stars with other different properties. We find that various possible composition of compact stars substantially affects the oscillation frequencies. However, the mass-scaled angular $f$ mode frequency ($\omega_{f}M$) varies universally with compactness ($C$) for all hadronic, quark and hybrid stars. The $f$ mode frequency ($f_{f_{1.4}}$) of the canonical 1.4 $M_{\odot}$ compact star, obtained with different composition, is quite correlated with the canonical radius ($R_{1.4}$) and tidal deformability ($\Lambda_{1.4}$) while $f_{p_{1.4}}$ is well correlated with slope parameter of the symmetry energy. We also show that $f_{g_{1.4}}$ of the HSs varies almost linearly with $\Lambda_{1.4}$. Should $g$ modes be detected, they could not only support the existence of HSs, but $f_{g}$ could be useful to understand the strength of quark repulsion in HSs.
@@ I Introduction
Several extreme conditions of density ($\rho$ = 5 - 10 times nuclear matter density), compactness ($C$ = $M/R$), and etc. related to the neutron star (NS) environment make them one of the most fascinating celestial objects to study. However, till date a lot of uncertainties still pertain to their equation of state (EoS), which is developed on the basis of theoretical modeling of NS matter. These uncertainties are generally attributed to the fact that the composition, properties and interaction of matter at such high density are largely inconclusive. Over several decades theoretical speculations suggest that NS cores are dense enough to sustain the existence of exotic matter and even signatures of hadron-quark phase transition thereby forming hybrid stars (HSs). In order to achieve phase transition, mechanisms like Gibbs and Maxwell constructions Maruyama et al. (2008); Orsaria et al. (2019), hadron-quark crossover Qin et al. (2023); Constantinou et al. (2021); Sotani and Kojo (2023) and constant speed of sound parameterization Sun and Wen (2023); Laskos-Patkos and Moustakidis (2023) etc. are often adopted in literature. The Bodmer-Witten conjecture suggests that strange quark matter (SQM), being more stable than pure nucleonic matter, is the true ground state of the hadronic matter Farhi and Jaffe (1984); Torres and Menezes (2013); Olinto (1987). This implies that the entire compact star may be composed of SQM and thereby strange quark stars (SQSs) may also exist. At very high density, QM can also lead to the formation of color-flavor-locked (CFL) QM Rajagopal and Wilczek (2001); Alford and Reddy (2003); Lugones and Horvath (2002); Vásquez Flores and Lugones (2017). Consequently, on the basis of the composition and interactions, we obtain the EoS of hadronic neutron stars (HNSs), SQSs, and HSs. Each class of EoS has its own uncertainty revolving around the consideration of composition and interactions. With the recent developments in the observational sector, the EoS has been constrained to a certain extent. The most important of such observational constraints include the detection of gravitational wave GW170817 by the LIGO/Vigro Collaboration Abbott et al. (2018) and the simultaneous measurement of mass and radius of HESS J1731-347 V. Doroshenko, V. Suleimanov, G. Pühlhofer, and A. Santangelo (2022) and those of PSR J0030+0451 Riley et al. (2019); Miller et al. (2019) and PSR J0740+6620 Fonseca et al. (2021); Miller et al. (2021); Riley et al. (2021) by NICER experiment.
@@
With the recent advancements of the upcoming gravitational wave (GW) detectors such as the LIGO O4 run, the Einstein Telescope Punturo et al. (2010); Hild et al. (2011), and the Cosmic Explorer Reitze et al. (2019); Evans et al. (2023), it can be expected that the oscillation mode frequencies of the GWs can be detected. This will provide deeper insight into the interior of the compact stars. Such oscillation spectra of GWs consist of different modes viz., fundamental ($f$), pressure ($p$), rotational ($r$), space-time ($w$) and gravity ($g$) modes Kokkotas and Schmidt (1999). The restoring force that re-establishes the equilibrium is different for each mode. In case of both the $f$ and $p$ modes, the equilibrium is restored by the fluid pressure of the star, and for the $g$ mode, gravity acts as the restoring force. The efficiency of the upcoming detectors is expected to help the branch of GW asteroseismology, especially in terms of the most prominent $f$ mode frequency Abac et al. (2025). This is because compared to the other modes, $f$ mode has the strongest tidal coupling and below resonance the behavior of $f$ mode is the primary contributor to the tidal deformability. Therefore, the detection and measurement of the $f$ mode frequency is the most plausible in the upcoming detectors. The other modes, especially the $g$ mode, are comparatively difficult to detect because their tidal couplings are much weaker. Considering the advances in the upcoming detectors, it is important to compute the oscillation mode frequencies with the different EoS. Several studies have been conducted on these different oscillation modes Sotani et al. (2011); Flores and Lugones (2014); Andersson and Pnigouras (2019); Zhao and Lattimer (2022); Kumar et al. (2024a); Ranea-Sandoval et al. (2018); Vásquez Flores et al. (2019); Pradhan et al. (2022); Jyothilakshmi et al. (2024a); Das et al. (2021); Hong et al. (2023); Shirke et al. (2024); Jaiswal and Chatterjee (2021); Constantinou et al. (2021); Orsaria et al. (2019); Kumar et al. (2023a); Zheng et al. (2023); Sen and Guha (2024); Wen et al. (2019); Sotani (2021); Jyothilakshmi et al. (2024b); Thakur et al. (2024a) based on different compositions of the star. Several works dedicated to this topic suggest strong correlation between the $f_{f}$ and compactness Andersson and Kokkotas (1998), moment of inertia Lau et al. (2010) and static tidal polarizability Chan et al. (2014); Sotani and Kumar (2021) for NSs, QSs and HSs Zhao and Lattimer (2022). This leads to the emergence of the universal relations between such properties of compact stars with $f_{f}$. $p$ mode frequency ($f_{p1}$), on the other hand, are weakly correlated with such properties of NSs Andersson and Kokkotas (1998). Moreover, they are largely affected by the EoS of the crust, unlike the $f_{f}$, especially at low density Vásquez Flores et al. (2017); Pradhan and Chatterjee (2021); Kunjipurayil et al. (2022). At zero temperature ($T$=0) the $g$ mode and its corresponding frequency $f_{g}$ arise due to the discontinuity in density of the NS, which may be attributed to the phenomena like first order phase transition. Thus it can be described by HS EoS Ranea-Sandoval et al. (2019); Flores and Lugones (2014); Sotani et al. (2011); Ranea-Sandoval et al. (2018); Tonetto and Lugones (2020); Wei et al. (2020); Rodriguez et al. (2021); Kumar et al. (2023b); Zhao et al. (2022); Tran et al. (2023); Zhang et al. (2024). Universal relations are found to exist between the mass-scaled $g_{1}$ mode frequency and the radius-scaled $f$ mode frequency; and also between the compactness-scaled $g_{1}$ mode frequency and tidal deformability Kuan et al. (2022). For an isolated, self-bound star, the $g$ mode has weakest tidal coupling and $g$ mode emission in the GWs is much weaker than the preceding modes. Therefore, it becomes quite challenging to detect $g$ mode in near future. However, in a special case of binary merger, that involves at least one HS, the tidal energy of the resonant $g$ modes can be twice or thrice the same for normal NS. This can lead to considerable excitation of the $g$ mode that may aid the detection of $f_{g}$ in future Jaikumar et al. (2021).
@@
In the present work, we study the effects of the different composition of compact star matter on the non-radial oscillation frequencies of the $f$ and $p$ modes of the HNSs and SQSs with different models. For HSs we also calculate the $g$ mode frequency $f_{g}$. The first development of the theory of non-radial oscillation of NS in the framework of general relativity (GR) is done in Thorne and Campolattaro (1967) while Lindblom and Detweiler (1983) first integrated the numerical solution of the NS oscillation. Later, the metric perturbations were neglected with the help of Cowling approximations Cowling (1941) in order to obtain a simplified theory Sotani et al. (2011). The calculation of the oscillation frequencies in full GR conditions that include the spacetime oscillation also shows the emergence of another mode viz., the $w$ mode. Therefore, many recent works have calculated the oscillation frequencies in the framework of the total GR theory Shirke et al. (2024); Mondal and Bagchi (2024); Zhao and Lattimer (2022); Kunjipurayil et al. (2022); Pradhan et al. (2022); Flores et al. (2024); Tsui and Leung (2005). The results of the calculation of the oscillation frequencies using Cowling approximation and total GR calculations can differ up to 30% for the $f_{f}$, and $\sim$15for $f_{p1}$ Kunjipurayil et al. (2022). However, the overestimation with Cowling approximation can be less than 20% from those calculated by considering the full GR treatment Pradhan et al. (2022); Yoshida and Kojima (1997). Also, Tonetto and Lugones (2020); Zhao et al. (2022) showed that the Cowling approximation is reasonably good for the calculation of $f_{g}$. Overall, the gross qualitative results remain unaffected with or without Cowling approximation. Therefore, it is still widely adopted in the recent literature Kumar et al. (2023b); Jyothilakshmi et al. (2024a); Arbañil et al. (2023); Thakur et al. (2024b); Thapa et al. (2023). In the present work we also consider the Cowling approximation in order to provide the estimates of different oscillation frequencies. For the purpose of exploring the uncertainty due to the EoS, we consider the KIDS (Korea-IBS-Daegu-SKKU) functional Gil et al. (2021); Gil and Hyun (2021); Gil et al. (2022), and the relativistic mean field (RMF) models GM1 Glendenning and Moszkowski (1991) and DD2 Typel et al. (2010) hadronic models. For SQSs we consider the vector Bag (v-Bag) model Torres and Menezes (2013); Lopes et al. (2022), the Nambu-Jona-Lasino (NJL) model with vector interaction Lopes and Menezes (2021); Masuda et al. (2013), and the CFL model based on the MIT Bag model Lugones and Horvath (2002). For HSs, we follow Maxwell construction Kumar et al. (2023c); Qin et al. (2023); Laskos-Patkos et al. (2024) with the KIDS model to describe the hadronic phase and the v-Bag model for the quark phase. We intend to study how the oscillation properties vary for the three different types of stars viz. HNSs, SQSs and HSs based on the composition. In this context, we mention that the KIDS model is not considered before in literature to study the oscillation properties of HNSs and HSs.
@@
We organize this paper in the following manner. In the next Sec. II we discuss the salient features of the different models considered in the present work on the basis of the different composition. Sec. III is devoted to the description of hadron-quark phase transition. We also provide the essentials of the mechanism to obtain the structural properties and different modes of non-radial oscillation of the compact star in Sec. IV. We display our results with corresponding discussion in Sec. V. We finally conclude in Sec. VI.
@@ II Models for hadronic and quark matters
We consider the KIDS functional Gil et al. (2021); Gil and Hyun (2021); Gil et al. (2022) as the hadronic model to compute the structural properties and the oscillations frequencies of HNSs. We compare the results obtained with the KIDS model, with those of the two RMF models viz., the GM1 Glendenning and Moszkowski (1991) and the DD2 Typel et al. (2010) models.
The KIDS models are selected to have the ranges of the symmetry energy parameters that are consistent with both nuclear data and astronomical observation of the neutron stars. Symmetry energy parameters are defined by
$S(\rho)=J+Lx+\frac{1}{2}K_{\rm sym}x^{2}+\cdots$ (1)
where $S$ denotes the symmetry energy, $x=(\rho-\rho_{0})/3\rho_{0}$, $\rho$ is the matter density and $\rho_{0}$ is the nuclear saturation density. The parameter most relevant to the NS properties is the slope parameter $L$, and the range obtained from the KIDS model is 47–66 MeV. This is consistent with empirical range 40–76 MeV Dutra et al. (2012), but the result of PREX-2 indicates that $L$ could be larger 80 MeV Reed et al. (2021). The possibility of $L$ above the empirical range is accounted for by considering two well-known RMF models viz., GM1 and DD2. Among them, GM1 has density-independent and DD2 has density-dependent nucleon-meson couplings. The two models are selected to have both similarity and difference with the KIDS models. In the DD2 model, the value of $L$ is within the empirical range, so it is similar to the KIDS model. However, the sign of $K_{\rm sym}$ is negative in the KIDS model, but positive in the DD2 model. Since $K_{\rm sym}$ also plays a critical role in the EoS of neutron star matter, the DD2 model can produce EoS stiffer than the KIDS model. Another RMF model GM1 has $L=94$ MeV, so one can easily expect that the model will give EoS much stiffer than the KIDS models. Thus, considering the four KIDS and two RMF models, the variation of $L$ ranges from 47 MeV (KIDS-D) to 94 MeV (GM1). The result, as will be shown soon, turns out as expected, so we can explore the impact of the stiffness of EoS to the oscillation mode.
To obtain the structural properties and the oscillations frequencies of the QSs we choose to work with different quark matter models viz., the vector Bag (v-Bag) model Torres and Menezes (2013); Lopes et al. (2022), the Nambu-Jona-Lasino (NJL) model with vector interaction Lopes and Menezes (2021); Masuda et al. (2013), and the color-flavor-locked (CFL) model based on the MIT Bag model Lugones and Horvath (2002).
All three quark models are quite well-known and have both similarities and differences among themselves. The unpaired quark models viz. the v-Bag and the NJL model have similarities that both the models consider the effect of vector interaction between quarks. However, the treatments of the two models are very different. The v-Bag model is based on the notion of confinement of the quarks with hypothetical ‘bag’ while the NJL model considers the four-Fermi contact type of interaction. The CFL phase is also important at high density, which also introduces the role of the pairing gap. Therefore, the choice of the hadronic and quark models is based on different treatments. Therefore, we consider a variety of models in the context of treatment and properties to study their effects on the structural and oscillation properties. The salient features of each of the hadronic and quark models, considered in the present work, are briefly addressed below.